# Synthetic Data Generation Using Large Language Models: Advances in Text and Code

### v1

@@ Abstract
Large language models (LLMs) have unlocked new possibilities for generating synthetic training data in both natural language and code domains. By producing artificial but task-relevant examples, these models can significantly augment or even substitute for real-world datasets, particularly in scenarios where labeled data is scarce, expensive, or sensitive. This paper surveys recent advances in leveraging LLMs to create synthetic text and code, highlighting key techniques such as prompt-based generation, retrieval-augmented pipelines, and iterative self-refinement. We examine how these methods can enrich low-resource tasks (e.g. classification, question answering) and facilitate code-centric applications (e.g. instruction tuning, code translation, bug repair) through automated verification of functional correctness. Alongside potential benefits—cost-effectiveness, broad coverage, and controllable diversity—we discuss the accompanying challenges, including factual inaccuracies in generated text, insufficient stylistic or distributional realism, and risks of bias amplification. Proposed mitigation strategies range from filtering and weighting synthetic outputs to reinforcement learning with execution feedback in code domains. We conclude by outlining open research directions, such as automated prompt engineering, cross-modal data synthesis, and robust evaluation frameworks, underscoring the growing importance of LLM-generated synthetic data in accelerating AI development while emphasizing ethical and quality safeguards.
@@ 1 Introduction
  • Challenges and Mitigations: We outline the main challenges in using LLMs for synthetic data, such as ensuring data fidelity and factuality, avoiding modeling biases, maintaining diversity without drifting from real data distribution, and preventing model collapse from iterative self-training on AI-generated data. We discuss solutions proposed in literature, like filtering and weighting synthetic examples [33], blending synthetic and real data [18], and leveraging feedback (e.g., executors for code, retrieval for facts) to improve quality [17] [28].
@@ 3.1 Various approaches for Prompt-Based Synthetic Text Data Generation
Prompt-based generation has been empirically shown to produce useful training data. A case study by Zhuoyan Li et al. [35] used GPT-3.5 to generate text classification data under various prompting methods. They evaluated the quality by training classifiers on the synthetic data and testing on real validation sets. A key finding was that in low-resource settings, synthetic data significantly boosts performance. Specifically, when only 100 real training examples were available, augmenting with 100 additional GPT-generated examples yielded 3%–26% absolute improvement in accuracy/F1 across several classification tasks. This gain was most pronounced in tasks where the model was severely underfit due to data paucity (e.g., a 26% jump in a news classification task with few real samples) [3]. However, as the amount of real data increases, the marginal benefit of synthetic data diminishes. With 1000 real examples, Yu et al. [60] observed only minor improvements (<5%) from adding synthetic data, and in some tasks no improvement. This suggests that synthetic augmentation is most useful for low-data regimes, and that purely synthetic data still has a quality gap compared to large real datasets [60]. In fact, prior works noted that models trained on an equal number of synthetic vs. real examples usually perform worse with the synthetic ones [60, 51]. Thus, synthetic data is better viewed as a supplement to small real datasets, rather than a wholesale replacement for large real corpora (except in zero-data scenarios) [39, 61].
Another insight from prompt-based generation studies is the trade-off between diversity and relevance. Zero-shot prompting can produce a wide range of outputs if the model isn’t constrained by examples, but some outputs may be off-target or less relevant [11]. Few-shot prompting, on the other hand, produces on-target outputs but risks many duplicates or very similar phrasing since the model mimics the few given examples [24]. Yu et al. [39] found that the zero-shot topic method, which explicitly varied the prompt for each generation via different topics, yielded the most diverse synthetic datasets and often the best downstream performance in the synthetic-only setting [27]. Datasets with higher diversity (covering varied subtopics) tended to train more robust models [62]. Conversely, a plain zero-shot method (same prompt each time) gave the least diverse data and worst performance on several tasks [62]. Interestingly, when synthetic data was mixed with real data, the importance of maximizing diversity dropped, as the real data provided a baseline of variation. In such cases, even one-shot or few-shot generated data (which may be conservative in diversity) helped by aligning well with real examples [62]. The best practice emerging is to use prompt designs that encourage diversity (via randomization or controlled variation) especially if relying heavily on synthetic data, while ensuring the generated examples remain realistic and on-task (using in-context examples or instructions to ground the model) [15].
Advanced Prompting Strategies Researchers have also developed iterative prompting strategies to refine synthetic data quality. For instance, Self-Instruct [53] is an approach where after an initial round of generation, one identifies difficult cases where a model still makes mistakes (misclassified examples) and then iteratively prompts the LLM to generate more data similar to those hard cases. By focusing on the model’s errors, this method aims to reduce the gap between the synthetic data distribution and the true data distribution, essentially targeting the weaknesses of the model. Such feedback loops can yield synthetic data that incrementally pushes model performance on challenging aspects of the task. Another approach, SunGen [16], acknowledges that not all LLM outputs are perfect and some synthetic examples may be noisy or low-quality. SunGen applies a weighted loss during model training, giving lower weight to potentially noisy synthetic examples so they do not harm the model [16]. In essence, it learns with a down-weighted penalty on data suspected to be hallucinated or label-noisy. This kind of selective training can make the model more robust to imperfect synthetic data.
@@
Table LABEL:tab:synthetic_data_methods summarizes a few representative methods and studies in LLM-based text data generation, highlighting their target tasks and key contributions.
Table 2: Selected LLM-based synthetic text data generation approaches and their key features.
@@ 3.2 Quality and properties of generated data
Synthetic QA and Reasoning An emerging area is using LLMs to generate not just answers but also explanations or reasoning chains for QA tasks. For instance, to train models that perform multi-hop reasoning, one can prompt an LLM to generate a question, the correct answer, and a step-by-step explanation (chain-of-thought). This produces synthetic training data for both the task and the format of reasoning. However, ensuring the correctness of the chain-of-thought is challenging; often an LLM might reach a correct answer via a flawed explanation, or vice versa. Therefore, researchers sometimes constrain generation by verifying final answers or using tools (like calculators, knowledge bases) in the loop to confirm intermediate facts.
@@ 4.1 Various approaches for Prompt-Based Synthetic Code Generation
Large language models capable of coding (e.g., OpenAI Codex, CodeGen, Code-Llama, StarCoder) are typically trained on massive corpora of existing code from repositories. However, researchers have started to supplement real code corpora with synthetic code generated by LLMs to address specific gaps or to fine-tune models on particular tasks [6]. Synthetic code data generation can take multiple forms, as creation of intruction-based datasets, code translation and refactoring, generating problem’s specifications and solutions.
@@ 4.1.3 Problem Generation and Solution Synthesis
Table 3: Selected LLM-based synthetic code data generation approaches.
Begin of Table
@@
Approach / System Code Task Domain Key Idea and Contribution Input Type Benchmark/Dataset Used
@@
CodeRL [31] Generation of problem’s specification and solution – Code generation with RL Uses an actor-critic reinforcement learning setup on code generation: the LLM generates code, an execution feedback (did the code run correctly?) serves as a reward, and the model is refined with this feedback. Code2Code HumanEval[5], MultiPL-E[21]
@@
Self-Code-Improve [19] Code translation and refactoring – Code optimization / repair LLM generates code and its own chain-of-thought for solving a task, then analyzes the code for errors or inefficiencies using heuristic prompts (simulated self-debugging). Code2Code No specific benchmark mentioned, but aligns with execution-based debugging tasks.
@@
InterCode [57] Generation of problem’s specification and solution – Interactive code generation (RL) Framework combining LLMs with an interactive environment: the LLM produces code as actions and receives execution feedback as observations. Code2Code No specific benchmark mentioned, aligns with execution-based evaluations.
@@
Reflexion [49] Generation of problem’s specification and solution – Self-refinement (general) A strategy in which an LLM first outputs a solution, then refines it based on feedback, creating pairs of (initial attempt, refined attempt). Code2Code HumanEval[5]
@@
Code Alpaca [4] Creation of instruction-based datasets – Code instruction following A 20K dataset of synthetic coding instructions and answers generated via Self-Instruct using ChatGPT. Text2Code No specific benchmark, aligns with instruction-based evaluations.
@@
WizardCoder [40] Creation of instruction-based datasets – Complex code instructions Built on Code Alpaca, WizardCoder applied an evolutionary approach to generate more complex and diverse instructions. Text2Code HumanEval[5], CoderEval[59]
@@
Magicoder – OSS-Instruct [55] Creation of instruction-based datasets – Code instruction tuning Created 75K synthetic Q&A pairs from open-source code. Text2Code No specific benchmark, aligns with instruction-tuned datasets.
@@
Test-Translating for Low-Resource [2] Code translation and refactoring – Multilingual coding Proposed generating code in low-resource programming languages by translating high-resource code and verifying correctness. Code2Code No specific benchmark mentioned, aligns with multilingual dataset evaluations.
@@
NL2SQL Augmentation [38] Creation of instruction-based datasets – Text-to-SQL parsing Combines strong and weak LLMs to generate diverse SQL queries. Text2Code No specific benchmark mentioned, aligns with text-to-SQL evaluation.
@@
PanGu-Coder [7] Creation of instruction-based datasets – Code generation A large-scale LLM trained by Huawei for code generation. Text2Code No specific benchmark mentioned.
@@
CodeGen [44] Creation of instruction-based datasets – Code generation and completion Open-source model designed for text-to-code generation across multiple programming languages. Text2Code HumanEval[5], MultiPL-E[21]
@@
InCoder [14] Creation of instruction-based datasets – Code completion and synthesis Generative model for code completion, predicting and generating code snippets autoregressively. Text2Code HumanEval[5], DS-1000[30]
@@
AlphaCode [34] Generation of problem’s specification and solution – Competitive programming Developed by Microsoft, AlphaCode generates multiple solutions per problem and ranks them by correctness. Text2Code HumanEval[5], CoderEval[59]
@@
Codex [5] Creation of instruction-based datasets – General code generation OpenAI’s Codex powers GitHub Copilot, pretraining on both natural language and source code. Text2Code HumanEval[5], MultiPL-E[21]
@@
TransCoder [29] Code translation and refactoring – Code translation An unsupervised machine translation model for programming languages. Code2Code No specific benchmark, aligns with multilingual model evaluations.
@@
LintSeq [46] Code translation and refactoring – Incremental code modifications Generates synthetic data by taking real code and creating incremental edits. Code2Code No specific benchmark, aligns with execution and edit-based dataset evaluations.
@@
AlphaDev [41] Generation of problem’s specification and solution – Optimized code synthesis DeepMind’s reinforcement learning-based model for discovering more efficient sorting algorithms. Code2Code AiXBench[20]
@@
LLM-itation (Buggy Code Generation) [32] Creation of instruction-based datasets – Buggy code and fix generation Uses LLMs (GPT-4) to generate synthetic buggy student code submissions for tutoring. Text2Code No specific benchmark, aligns with bug-detection models and educational datasets.

### v2

@@ Abstract
This survey reviews how large language models (LLMs) are transforming synthetic training data generation in both natural language and code domains. By producing artificial but task-relevant examples, these models can significantly augment or even substitute for real-world datasets, particularly in scenarios where labeled data is scarce, expensive, or sensitive. This paper surveys recent advances in leveraging LLMs to create synthetic text and code, highlighting key techniques such as prompt-based generation, retrieval-augmented pipelines, and iterative self-refinement. We examine how these methods can enrich low-resource tasks (e.g., classification, question answering) and facilitate code-centric applications (e.g., instruction tuning, code translation, bug repair) through automated verification of functional correctness. Alongside potential benefits—cost-effectiveness, broad coverage, and controllable diversity—we discuss the accompanying challenges, including factual inaccuracies in generated text, insufficient stylistic or distributional realism, and risks of bias amplification. Proposed mitigation strategies range from filtering and weighting synthetic outputs to reinforcement learning with execution feedback in code domains. We conclude by outlining open research directions, such as automated prompt engineering, cross-modal data synthesis, and robust evaluation frameworks, underscoring the growing importance of LLM-generated synthetic data in accelerating AI development while emphasizing ethical and quality safeguards.
Keywords Synthetic Data Generation  $\cdot$ Large Language Models (LLMs)  $\cdot$ Text Data Augmentation  $\cdot$ Code Data Synthesis  $\cdot$ Prompt Engineering  $\cdot$ Instruction Tuning  $\cdot$ Machine Learning Training Data  $\cdot$ Natural Language Processing (NLP)  $\cdot$ Code Generation  $\cdot$ Reinforcement Learning for Code  $\cdot$ Automated Data Annotation  $\cdot$ Bias and Fairness in Synthetic Data  $\cdot$ Retrieval-Augmented Generation (RAG)  $\cdot$ Evaluation of Synthetic Data  $\cdot$ Model Collapse in LLMs
@@ 1 Introduction
  • Challenges and Mitigations: We outline the main challenges in using LLMs for synthetic data, such as ensuring data fidelity and factuality, avoiding modeling biases, maintaining diversity without drifting from real data distribution, and preventing model collapse  [^1] Model collapse refers to the degradation of model quality that occurs when successive generations of models are trained primarily or exclusively on synthetic data produced by other models, leading to loss of diversity, factuality, or robustness; see [18]. from iterative self-training on AI-generated data. We discuss solutions proposed in literature, like filtering and weighting synthetic examples [33], blending synthetic and real data [18], and leveraging feedback (e.g., executors for code, retrieval for facts) to improve quality [17] [28].
@@
To guide readers through the wide landscape of LLM-driven synthetic data generation, Table 1 presents a high-level taxonomy of tasks addressed in the literature. We distinguish between applications in the text and code domains, further categorizing major synthetic data tasks such as generation, translation, instruction tuning, and repair. This organizational scheme forms the backbone of our survey and is reflected in the structure of the following sections.
Table 1: Taxonomy of LLM-based synthetic data tasks
Text Domain Code Domain
Generation (augmentation, prompts) Generation (code synthesis)
Translation (cross-lingual) Translation (code-to-code)
Instruction Tuning (Alpaca, etc.) Instruction Tuning (Code Alpaca, WizardCoder)
Paraphrasing / Style Transfer Refactoring / Style Transfer
QA / Reasoning (chain-of-thought) QA / Problem Generation
Retrieval-Augmented Generation Retrieval-Augmented Generation
Evaluation & Filtering Evaluation (execution feedback)
Repair (N/A) Repair (bug fixing, code improvement)
@@
2 Related Work
The field of synthetic data generation, especially with large language models (LLMs), has evolved rapidly, prompting the publication of several recent survey articles that address various aspects of this topic. We briefly review and position our work with respect to these prior surveys.
Synthetic Data for NLP. Wang et al. [52] provide a comprehensive overview of data synthesis and augmentation techniques for large language models in natural language processing (NLP). Their survey covers a broad spectrum of approaches, from traditional text augmentation methods to modern neural and prompt-based strategies. However, their primary focus is on NLP applications, with only limited discussion of synthetic data for programming code or the distinctive challenges faced in code generation domains.
LLM-Driven Synthetic Data Pipelines. Long et al. [39] propose a generic framework for LLM-driven synthetic data generation, curation, and evaluation. Their review encompasses text, code, and other modalities, organizing the landscape around lifecycle stages such as data generation, filtering, and integration. While thorough in its coverage, the survey remains high-level, and does not systematically analyze or compare empirical findings for text and code tasks separately. Key challenges unique to code, such as leveraging execution feedback  [^2] Execution feedback uses the outcome of running generated code—such as passing or failing test cases—as a supervisory signal for selecting or further training models [34]. and handling functional correctness, are not addressed in detail.
Surveys on Synthetic Code Data. A number of surveys focus specifically on code generation with LLMs. Jiang et al. [26] review architectural advances and applications of large language models for code, with a primary emphasis on model capabilities and benchmark results. Chen et al. [6] present a rich taxonomy of synthetic data generation techniques for code LLMs, cataloging a wide array of recent approaches for code instruction tuning, translation, and program repair. While these works offer valuable overviews, they are largely centered on code, with less attention given to synthetic data issues shared across natural language and code domains, such as distributional alignment, bias amplification, or unified evaluation strategies.
Position and Novelty of This Survey. In contrast to these prior works, our survey aims to bridge the gap between synthetic data generation for natural language and programming code, providing a unified and comparative treatment of LLM-driven methods across both modalities. Specifically, our contributions are:
  • Unified Framework: We synthesize and extend recent advances in both text and code synthetic data generation, providing a cross-domain taxonomy of LLM-driven techniques—including prompt-based, retrieval-augmented, and iterative self-refinement pipelines.
  • Comparative Analysis: We systematically compare empirical results, challenges, and mitigation strategies in both domains, with particular focus on issues of quality assurance, distributional realism, bias, and model collapse, which are often addressed separately in prior surveys.
  • Practical Lessons and Future Directions: We distill practical recommendations from recent studies and outline a forward-looking research agenda that incorporates trends in evaluation, human-in-the-loop pipelines, and ethical considerations relevant to both NLP and code generation communities.
Thus, this survey advances the literature by offering greater scope—through the joint treatment of text and code tasks—greater depth via comparative empirical analysis, and a novel structure that unifies recent methods, challenges, and practical insights across modalities.
3 Survey Methodology
To ensure transparency and reproducibility, we detail here the systematic methodology employed for selecting and reviewing the literature included in this survey, following a PRISMA-style approach as summarized in Figure 1.
Inclusion Criteria: We considered papers and preprints that met the following criteria:
  • Published between January 2020 and April 2025, to capture the period of rapid advancement in LLM-driven synthetic data generation.
  • Address synthetic data generation, data augmentation, or instruction tuning in the context of large language models (LLMs), for either text or code domains.
  • Focus on models with at least several hundred million parameters (to ensure relevance to the LLM paradigm), including both proprietary (e.g., GPT-3/4, Claude, Gemini) and open-source (e.g., Llama, CodeLlama, StarCoder) models.
  • Cover a range of task categories, including but not limited to: classification, question answering, instruction following, code synthesis, code translation, and bug repair.
  • Provide empirical evidence, systematic evaluation, or substantial methodological innovation relevant to synthetic data workflows.
Exclusion Criteria:
  • Works focused exclusively on synthetic data for modalities outside text or code (e.g., images, speech) unless providing transferable methodology or insights.
  • Non-peer-reviewed content lacking methodological rigor or empirical validation (e.g., blog posts, opinion pieces).
  • Papers predating 2020 or focused on pre-LLM generative models, unless foundational to understanding current techniques (e.g., early data augmentation methods).
Review Process: The literature review process was conducted in a semi-automated manner. We performed systematic keyword searches on Google Scholar, arXiv, and Semantic Scholar using terms such as “synthetic data,” “LLM augmentation,” “instruction tuning,” “prompt engineering,” “code generation,” and related task names (e.g., “text classification,” “code translation”). Reference lists of recent surveys and benchmark papers were also examined to identify additional relevant studies (snowballing). Titles and abstracts were screened to exclude irrelevant works, and full texts were reviewed for eligibility based on the inclusion/exclusion criteria above. The final set of papers reflects a curated sample that spans both influential early works and state-of-the-art studies published as of April 2025.
Literature Screening Flow: To increase transparency, Figure 1 summarizes the identification, screening, and selection process for works included in this review, following PRISMA guidelines.
Figure 1: PRISMA-style flow diagram for study selection.
@@ 5.1 Various Approaches for Prompt-Based Synthetic Text Data Generation
Prompt-based synthetic data generation, discussed in detail in Section 4.2, forms the backbone of modern LLM-driven augmentation in NLP. As previously outlined, the choice of prompting strategy—zero-shot, few-shot, or topic-controlled—critically impacts both the diversity and relevance of synthetic examples, with each approach presenting distinct trade-offs.
Recent empirical studies have systematically benchmarked these strategies. For instance, Li et al. [35] showed that, in low-resource settings, augmenting 100 real training samples with 100 GPT-3.5-generated synthetic samples can yield 3–26% improvements in accuracy or F1 across a range of text classification tasks. The benefit was most pronounced when data scarcity led to severe model underfitting, as also observed by Chai et al. [3]. However, as the quantity of real data increases, the marginal advantage of synthetic augmentation diminishes, with some studies reporting negligible improvements when synthetic data is added to large real datasets [60]. This trend suggests that synthetic data is especially valuable as a supplement in low-data regimes, rather than a full replacement for extensive, high-quality real corpora [51, 39, 61].
Beyond basic prompting, diverse strategies have emerged to further improve data utility and model robustness:
  • Topic-controlled or randomized prompting (e.g., WANLI [36], AugGPT [9]) systematically varies prompt content to maximize diversity across subtopics, which correlates with improved generalization—especially when training solely on synthetic data [39, 62].
  • Few-shot and instruction-based prompting are preferred where format precision or specific label alignment are required, as in GPT3Mix [58] and Unnatural Instructions [23]. These approaches can yield highly relevant and accurate examples, even if diversity is more constrained.
  • Iterative and feedback-driven generation (e.g., Self-Instruct [53], SunGen [16]) enhance dataset quality by focusing generation on failure cases, hard examples, or model weaknesses, often using automated weighting or human-in-the-loop validation to filter noisy or low-quality outputs.
Studies also note that the importance of maximizing diversity decreases when synthetic data is mixed with even modest amounts of real data, since the real samples anchor the data distribution and provide natural variety [62]. Furthermore, recent works encourage adaptive or hybrid prompting, along with selective down-weighting of suspected noisy synthetic examples during model training [16], to further enhance robustness.
@@
Table LABEL:tab:synthetic_data_methods summarizes several representative prompt-based synthetic data generation methods, along with their target tasks and main contributions.
Table 3: Representative LLM-based synthetic text data generation methods and their core contributions. This table reveals the rapid evolution of strategies—from prompt engineering and task bootstrapping to diverse augmentation and self-instruction—demonstrating how LLMs now support cost-effective, scalable data creation across multiple NLP tasks. It also highlights common themes such as mixing synthetic and real data, leveraging iterative refinement, and targeting low-resource or complex tasks to maximize model improvements.
@@ 5.2 Quality and properties of generated data
Synthetic QA and Reasoning An emerging area is using LLMs to generate not just answers but also explanations or reasoning chains for QA tasks. For instance, to train models that perform multi-hop reasoning, one can prompt an LLM to generate a question, the correct answer, and a step-by-step explanation (chain-of-thought  [^3] Chain-of-thought (CoT) prompting guides models to produce intermediate reasoning steps or explanations before providing a final answer [13].). This produces synthetic training data for both the task and the format of reasoning. However, ensuring the correctness of the chain-of-thought is challenging; often an LLM might reach a correct answer via a flawed explanation, or vice versa. Therefore, researchers sometimes constrain generation by verifying final answers or using tools (like calculators, knowledge bases) in the loop to confirm intermediate facts.
@@ 6.1 Various approaches for Prompt-Based Synthetic Code Generation
Large language models capable of coding (e.g., OpenAI Codex, CodeGen, Code-Llama, StarCoder) are typically trained on massive corpora of existing code from repositories. However, researchers have started to supplement real code corpora with synthetic code generated by LLMs to address specific gaps or to fine-tune models on particular tasks [6]. Synthetic code data generation can take multiple forms, as creation of instruction-based datasets, code translation and refactoring, generating problem specifications and solutions.
@@ 6.1.3 Problem Generation and Solution Synthesis
Table 4: Overview of representative LLM-based approaches for synthetic code data generation, covering tasks such as code generation, translation, refactoring, instruction tuning, and bug synthesis.
Note: Where “No specific benchmark” is listed, the method was evaluated on custom or task-specific datasets rather than a standard public benchmark.
@@
Approach / System Code Task Domain, Input Type Key Idea, Contribution, and Benchmark/Dataset
@@
CodeRL [31] Code generation with RL (Generation of problem specification and solution); Input: Code2Code Uses actor-critic RL on code generation. The LLM generates code; execution feedback (does the code run?) is used as a reward to refine the model. Benchmarks: HumanEval[5], MultiPL-E[21]
@@
Self-Code-Improve [19] Code translation, refactoring, optimization, repair; Input: Code2Code LLM generates code and its chain-of-thought, then analyzes for errors or inefficiencies (simulated self-debugging with heuristic prompts). Benchmark: No specific benchmark; aligns with execution-based debugging tasks.
@@
InterCode [57] Interactive code generation with RL (specification + solution); Input: Code2Code Combines LLMs with an interactive environment: code as actions, execution feedback as observations. Benchmark: No specific benchmark; execution-based evaluation.
@@
Reflexion [49] Self-refinement in code generation; Input: Code2Code LLM outputs an initial solution, then refines it based on feedback, creating (initial, refined) pairs. Benchmark: HumanEval[5]
@@
Code Alpaca [4] Code instruction following (instruction dataset creation); Input: Text2Code 20K dataset of synthetic coding instructions and answers via Self-Instruct with ChatGPT. Benchmark: No specific benchmark; instruction-based evaluations.
@@
WizardCoder [40] Complex code instructions (instruction dataset creation); Input: Text2Code Builds on Code Alpaca, using evolutionary approaches for more complex, diverse instructions. Benchmarks: HumanEval[5], CoderEval[59]
@@
Magicoder – OSS-Instruct [55] Code instruction tuning; Input: Text2Code Created 75K synthetic Q&A pairs from open-source code. Benchmark: No specific benchmark; aligns with instruction-tuned datasets.
@@
Test-Translating for Low-Resource [2] Multilingual code generation, translation; Input: Code2Code Generates code for low-resource languages by translating and verifying high-resource code. Benchmark: No specific benchmark; multilingual evaluations.
@@
NL2SQL Augmentation [38] Text-to-SQL parsing (instruction dataset creation); Input: Text2Code Combines strong and weak LLMs to generate diverse SQL queries. Benchmark: No specific benchmark; text-to-SQL evaluation.
@@
PanGu-Coder [7] Code generation (instruction dataset creation); Input: Text2Code Large-scale LLM for code generation (Huawei). Benchmark: No specific benchmark.
@@
CodeGen [44] Code generation and completion (instruction dataset creation); Input: Text2Code Open-source model for text-to-code generation in multiple languages. Benchmarks: HumanEval[5], MultiPL-E[21]
@@
InCoder [14] Code completion and synthesis (instruction dataset creation); Input: Text2Code Generative model for autoregressive code snippet completion. Benchmarks: HumanEval[5], DS-1000[30]
@@
AlphaCode [34] Competitive programming (problem spec & solution); Input: Text2Code Generates multiple solutions, ranks by correctness (Microsoft). Benchmarks: HumanEval[5], CoderEval[59]
@@
Codex [5] General code generation (instruction dataset creation); Input: Text2Code OpenAI’s Codex (GitHub Copilot), pretrained on both language and code. Benchmarks: HumanEval[5], MultiPL-E[21]
@@
TransCoder [29] Code translation (code2code); Input: Code2Code Unsupervised machine translation for programming languages. Benchmark: No specific benchmark; multilingual evaluation.
@@
LintSeq [46] Incremental code modifications (code2code); Input: Code2Code Generates synthetic data by creating incremental edits of real code. Benchmark: No specific benchmark; edit-based evaluation.
@@
AlphaDev [41] Optimized code synthesis (problem spec & solution); Input: Code2Code RL-based model for efficient sorting algorithms (DeepMind). Benchmark: AiXBench[20]
@@
LLM-itation (Buggy Code Generation) [32] Buggy code and fix generation (instruction dataset creation); Input: Text2Code Uses LLMs (GPT-4) to synthesize buggy student code for tutoring. Benchmark: No specific benchmark; educational and bug-detection datasets.